Suppose a wage increase from $12 to $16 an hour for Expedia call center reps increases the number of daily job applicants from 42 to 58. What is the price elasticity of labor supply?

Instructions: Round your response to one decimal place. Use the midpoint formula.
Answer #1

Answer

Elasticity of supply=(change in quantity/average quantity)/(change in price/average price)
change in quantity=58-42=16
average quantity=(58+42)/2=50
change in price=16-12=4
average price=(16+12)/2=14
Elasticity of supply=(16/50)/(4/14)

=1.12

the elasticity is 1.12

the supply is elastic as the elasticity is above 1.